OpenStudy (anonymous):

abc = 171 Solve for a,b,c

OpenStudy (y2o2):

a = 342/(2bc) b= 342/(2ac) c = 342/(2ab)

OpenStudy (y2o2):

a = 171/(bc) b= 171/(ac) c = 171/(ab)